Evaluate the ∫√5xdx


1
Expert's answer
2021-02-24T12:44:16-0500

The indefinite integral is evaluated as,


"\int\sqrt{5x}dx=\sqrt{5}\int\sqrt{x}dx"


Use formula, "\int x^ndx=\frac{x^{n+1}}{n+1}" to find the integral. Here, "n=\frac{1}{2}"


So, the integral is,


"\int\sqrt{5x}dx=\sqrt{5}[\frac{x^{\frac{1}{2}+1}}{\frac{1}{2}+1}]"


"=\sqrt{5}[\frac{x^{\frac{3}{2}}}{\frac{3}{2}}]"


"=\frac{2\sqrt{5}}{3}x^{\frac{3}{2}}"


Therefore, the indefinite integral is "\int\sqrt{5x}dx=\frac{2\sqrt{5}}{3}x^{\frac{3}{2}}"